OpenStudy (anonymous):

plss help!!!.....Determine if triangle DEF with coordinates D (2, 1), E (3, 5), and F (6, 2) is an equilateral triangle. Use evidence to support your claim. If it is not an equilateral triangle, what changes can be made to make it equilateral? Be specific.

OpenStudy (here_to_help15):

since u said plss i will help :p

OpenStudy (anonymous):

thanks soo much

OpenStudy (anonymous):

@Here_to_Help15

OpenStudy (here_to_help15):

yes

OpenStudy (here_to_help15):

im almost done waits ok

OpenStudy (anonymous):

kk @Here_to_Help15

OpenStudy (here_to_help15):

i dont think its a equilateral because d=squar root (x2-x1)^2+(y2-y1)^2 ab=(2,1) (3,5) squar root (3-2)^2+(5-1)^2= squar root 1+16= squar root 17 =4.123 bc=(3,5) (6,2) squar root (6-3)^2+(2-5)^2= squar root 9+9= squar root 18 = 4.242 ca=(6,2) (2,1) squar root (2-6)^2+(1-2)^2 squar root 16+1=squar root 17= 4.123

OpenStudy (here_to_help15):

@myrennazoe7 did that help at all?

OpenStudy (anonymous):

lol yeahh.. thanks @Here_to_Help15

OpenStudy (here_to_help15):

Yw :)